a girls age in 12 years time is the square of her present age. find her present age

Answer by Horlaye(27)

Let the present age be 'p'
In twelve years time
p + 12 = pē
pē - p - 12 = 0
pē + 3p - 4p - 12 = 0
p(p + 3) -4(p + 3) = 0
( p + 3 ) ( p - 4 ) = 0
p + 3 = 0 or p - 4 = 0
p = -3 or p = 4
Since one's age cannot be negative
Therefore her present age is p = 4